OpenStudy (anonymous):

Solve: 7(x − 1) = 5(x + 2) + 1 A=9 B=18 C=2 D=4

OpenStudy (aroub):

9

OpenStudy (anonymous):

7x - 7 = 5x + 10 + 1 7x - 7 = 5x + 11 2x = 18 x = 9 A, x = 9

OpenStudy (lalaly):

7x-7=5x+10+1 7x-5x=10+7+1 2x=18 x=9